OpenStudy (anonymous):

How do you know if a sequence is bounded( lower or upper bounded) ? Ex. An = n/(2^(n+2))

OpenStudy (anonymous):

Take the limit as x approaches infinity, nay?

OpenStudy (anonymous):

yes that is the right way

OpenStudy (anonymous):

f the limit exist the sequence is bounded

OpenStudy (kinggeorge):

Let me revise that statement I just made: If the sequence converges, it's bounded. If it diverges, it doesn't tell you anything.

OpenStudy (anonymous):

? Can't you say, if it diverges, that it's unbounded?

OpenStudy (anonymous):

no each convergence sequence is bounded and vice versa

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh wait wait. It could diverge but still be bounded.

OpenStudy (kinggeorge):

Take the sequence \[\sum_{n=0}^\infty ({-1})^n\]This is bounded but diverges.

OpenStudy (anonymous):

Wait, my book says that if a sequence is bounded and monotonic, it is convergent. My question is how to figure out if its bounded or not to find if its convergent?

OpenStudy (anonymous):

Right. Good example, Georgey.

OpenStudy (anonymous):

i think you are not right

OpenStudy (kinggeorge):

If it's monotonic and bounded, it is definitely convergent. My example is clearly not monotonic, which is why it's divergent.

OpenStudy (anonymous):

To prove it's bounded, you would need to find some thing (anything) that is greater than every term of the sequence.

OpenStudy (kinggeorge):

Or it could be bounded on the lower end, in which case just find something that's less than every term in the sequence.

OpenStudy (anonymous):

So...how do you find out where it is bounded at?

OpenStudy (anonymous):

For this sequence, notice that it's monotonically decreasing. Does it every reach 0?

OpenStudy (anonymous):

Pick 0 as a lower bound, and it's pretty easy to prove that any arbitrary term of the sequence is greater than 0.

OpenStudy (anonymous):

find the limit n tends to \[\infty\]

OpenStudy (kinggeorge):

If it's monotonic, just take the first term. It's either the lower bound or the upper bound depending on if the sequence is monotonically increasing or decreasing.

OpenStudy (kinggeorge):

Unless the first term is \(\pm \infty\). In that case, take the limit as \(n \rightarrow \infty\)

OpenStudy (anonymous):

Eh... the first term doesn't do it for you... if it's monotonically increasing, you need an upper bound to say that the limit exists. If it's monotonically decreasing, you need a lower bound to say that the limit exists. The first term will be a bound, but not the bound you need.

OpenStudy (anonymous):

hm..could you explain how i would find the bound for...say... (-1)^n(1/n)

OpenStudy (anonymous):

i think it is convergent to 0

OpenStudy (anonymous):

I think the best answer, in my experience is just to take some terms of the sequence and see if you can notice the pattern, what is happening with it. With the example you gave, I get A1 = 1/8 A2 = 2/16 =1/8 A3 = 3/32 A4 = 4/64 = 1/16 A5 = 5/128 A6 = 6/256 A7 = 7/512 As the denominator grows exponentially, the terms are getting very very small quickly. So I can see that the sequence decreases monotonically. If I can say that there's a lower bound that it never passes, then you can say that the limit exists. If you can say specifically what the GREATEST lower bound is, then THAT will be the limit. So, I notice it's getting quite small, but will never be negative, so the lower bound to pick becomes obvious. Pick 0, and prove that it's a lower bound.

OpenStudy (anonymous):

so as i said befor you find it convergent to 0

OpenStudy (anonymous):

For (-1)^n(1/n) A1 = -1 A2 = 1/2 A3 = -1/3 A4 = 1/4 A5 = -1/5 A6 = 1/6 and so on.